OpenStudy (anonymous):

Consider the polynomial x^3-3x-12. Show that this polynomial has a root between 1 and 2.

OpenStudy (aum):

f(x) = x^3 - 3x - 12 f(1) = ? f(2) = ? If f(x) changes sign between x = 1 and x = 2, then it must pass through zero somewhere between 1 and 2. So there will be a root between x = 1 and x = 2.

OpenStudy (aum):

Are you sure you typed the polynomial correctly? Because the one you typed does not have a root between 1 and 2.

OpenStudy (aum):

It has a root between 2 and 3.

OpenStudy (anonymous):

Sorry! It's x^3+3x-12 @aum

OpenStudy (aum):

(x) = x^3 + 3x - 12 f(1) = ? f(2) = ? If f(x) changes sign between x = 1 and x = 2, then it must pass through zero somewhere between 1 and 2. So there will be a root between x = 1 and x = 2 where f(x) = 0.